"An electrical firm manufactures light bulbs that have a length of life that is normally distributed with mean equal to 778 hours and a standard deviation of 43 hours. Find the probability that a bulb burns between 755 and 892 hours. 


1
Expert's answer
2021-03-01T15:34:37-0500

We have that

"\mu = 778"

"\sigma = 43"

"P(755<X<892) =P(\frac{755-\mu}{\sigma}<Z<\frac{892-\mu}{\sigma})=P(\frac{755-778}{43}<Z<\frac{892-778}{43})="

"=P(-0.53<Z<2.65)=P(Z<2.65)-P(Z<-0.53)="

"=0.9960-0.2981=0.6979"


Answer: 0.6979